### v1

# Entering Real Social World! Benchmarking the Theory of Mind and Socialization Capabilities of LLMs from a First-person Perspective

Abstract
In the social world, humans possess the capability to infer and reason about others’ mental states (such as emotions, beliefs, and intentions), known as the Theory of Mind (ToM). Simultaneously, humans’ own mental states evolve in response to social situations, a capability we refer to as socialization. Together, these capabilities form the foundation of human social interaction. In the era of artificial intelligence (AI), especially with the development of large language models (LLMs), we raise an intriguing question: How do LLMs perform in terms of ToM and socialization capabilities? And more broadly, can these AI models truly enter and navigate the real social world? Existing research evaluating LLMs’ ToM and socialization capabilities by positioning LLMs as passive observers from a third-person perspective, rather than as active participants. However, compared to the third-person perspective, observing and understanding the world from an ego-centric first-person perspective is a natural approach for both humans and AI agents. The ToM and socialization capabilities of LLMs from a first-person perspective, a crucial attribute for advancing embodied AI agents, remain unexplored. To answer the aforementioned questions and bridge the research gap, we introduce EgoSocialArena, a novel framework designed to evaluate and investigate the ToM and socialization capabilities of LLMs from a first-person perspective. It encompasses two evaluation environments: static environment and interactive environment, with seven scenarios: Daily Life, Counterfactual, New World, Blackjack, Number Guessing, and Limit Texas Hold’em, totaling 2,195 data entries. With EgoSocialArena, we have conducted a comprehensive evaluation of nine advanced LLMs and observed some key insights regarding the future development of LLMs as well as the capabilities levels of the most advanced LLMs currently available.
1 Introduction
In the complex social interactions of humans, two fundamental cognitive capabilities play crucial roles: Theory of Mind (ToM) and Socialization. ToM is a fundamental psychological process, defined as the capacity to reason about others’ mental states – beliefs, intents, desires, emotions, knowledge, etc (Premack & Woodruff, 1978; Ma et al., 2023; Chen et al., 2024). Socialization refers to the capability for own mental state evolution in response to social situations. As illustrated in Figure 1(A), When the little boy receives a birthday gift or achieves good grades, he feels very happy. These intertwined capabilities are central to humans’ social life.
With the advent of the era of LLMs, powerful models like GPT-4 (Achiam et al., 2023) and Claude (Anthropic, 2024) have demonstrated remarkable competence in multiple tasks and domains. It raises intriguing questions: How do LLMs perform in terms of ToM and socialization capabilities? And more broadly, can these AI models truly enter and navigate the real social world to achieve efficient human-computer interaction?
To evaluate the ToM and socialization capabilities of LLMs, multiple benchmarks have been proposed, such as SocialIQA (Sap et al., 2022), NormBank (Ziems et al., 2023), BigToM (Gandhi et al., 2023), FanToM (Kim et al., 2023), HI-ToM (Wu et al., 2023), OpenToM (Xu et al., 2024), NegotiationToM (Chan et al., 2024), and ToMBench (Chen et al., 2024). However, as illustrated in Figure 1(B), these benchmarks evaluate LLMs’ ToM and socialization capabilities by positioning LLMs as passive observers from a third-person perspective rather than as active participants. We propose two key points: (1) Observing and understanding the world from an ego-centric first-person perspective is a natural approach for both humans and AI agents. (2) Evaluating LLM’s ToM and socialization capabilities from an ego-centric first-person perspective measures whether LLMs can enter the real social world to achieve efficient human-computer interaction, providing indispensable resources for future research in the field of Embodied Artificial Intelligence (EAI).
Figure 1: (A): ToM (others mental states), Socialization (own mental states). (B): LLM acts as a passive observer to analyze the mental states of characters within a story from a third-person perspective. (C): Evaluation Scenarios in EgosocialArena. (D):Rule-based Agents and RL Agents within Interactive environment.
In this paper, we present EgosocialArena, a novel framework for evaluating and investigating LLMs’ ToM and socialization capabilities from a first-person perspective. We propose a structured taxonomy to guide the development of EgoSocialArena, encompassing three key components: (1) We propose a systematic methodology to transform two existing third-person ToM benchmarks, ToMI and ToMbench, into a first-person perspective. We present a comprehensive argument to justify the validity and rationale behind this transformation. The higher-order ToM questions in a third-person perspective are transformed to evaluate the first-person ToM capabilities of LLMs. In contrast, first-order ToM questions in a third-person perspective are transformed to evaluate LLMs’ mental states in response to social events, i.e., socialization capabilities. Detailed descriptions can be found in 3.1. (2) We investigate the socialization capabilities of LLMs in intriguing and distinctive social situations, including counterfactual, new world, and blackjack game scenarios, as illustrated in Figure 1(C). (3) We investigate the ToM capabilities of LLMs within interactive environments, including belief about the opponent’s behavior patterns in number-guessing scenarios and modeling of opponent’s play styles in Texas Hold’em poker game scenarios. In interactions between models of varying strengths, such as GPT-3.5 and GPT-4, we observe a ”babysitting” issue, where the weaker model can negatively impact the stronger one, diminishing its performance. To mitigate this issue and enable fair evaluation, as illustrated in Figure 1(D), we construct rule-based agents at different cognitive levels and train Reinforcement Learning (RL) agents to provide stable capabilities and behavioral strategies. Overall, EgoSocialArena encompasses two evaluation environments: static environment and interactive environment, with seven scenarios: Daily Life, Counterfactual, New World, Blackjack, Number Guessing, and Limit Texas Hold’em, totaling 2,195 data entries.
We conduct extensive experiments on EgoSocialArena to evaluate 9 foundational models known for their leading performance across multiple tasks and domains. This set includes five API-based models (i.e., o1-preview, GPT-4o, GPT-4-Turbo, GPT-3.5-Turbo, and claude-3-5-sonnet-20240620) and four open-source models (LLaMa-3-8B-Chat, LLaMa-3-70B-Chat, LLaMa-3-8B-Instruct, and LLaMa-3.1-405B-Instruct). We establish a human performance baseline by engaging qualified human annotators. Our experimental results reveal several interesting and critical insights:
  1. 1.
    Although LLMs currently perform behind humans in most scenarios, they have shown significant potential in a few specific cases.
  2. 2.
    The ToM capabilities of LLMs show significant differences between a third-person perspective and a first-person perspective.
  3. 3.
    The powerful capabilities of the o1-preview model are truly surprising.
  4. 4.
    The scaling up of open-source models or performing instruction fine-tuning has not yielded significant results in terms of ToM and socialization capabilities.
  5. 5.
    The performance gap between open-source models and API-based models in socialization scenarios is significant.
2 Related Work
Existing ToM Benchmarks
Previous evaluations for the ToM of LLMs primarily focus on testing models using narrative stories, also referred to as reading comprehension scenarios. Specifically, Le et al. (2019) proposed the ToMi benchmark based on the classic Sally-Anne test. Wu et al. (2023) introduced the HI-ToM benchmark, which focuses on higher-order belief reasoning and sets up scenarios where agents can communicate with each other. Gandhi et al. (2023) proposed BigToM, which presents a framework for designing a ToM benchmark from synthetic templates for evaluating different aspects of LLMs’ ToM capabilities (e.g., desire and belief). Xu et al. (2024) introduced OpenToM, which assigns personalities to agents in the stories and ensures that the storylines are more reasonable and logical. Chen et al. (2024) proposed ToMBench, which systematically evaluates LLMs across all dimensions of ToM capabilities. Unlike the above methods that require LLMs to read stories and answer related questions, some studies evaluate LLMs’ performance by inputting dialogues to them. Kim et al. (2023) proposed FanToM, which tests LLMs on their ability to infer the mental states of characters in everyday conversations. Chan et al. (2024) introduced NegotiationToM, which restricts the dialogue content to negotiation scenarios.
Existing Socialization Benchmarks
Sap et al. (2022) proposed SocialIQA and used it to evaluate LLMs. SocialIQA contains many questions related to social commonsense. Ziems et al. (2023) introduced NormBank, a large repository of social norms knowledge, which can be used to assess social norm-related tasks. Li et al. (2024) reorganized and classified existing datasets related to social intelligence. Xu et al. (2023) studied LLMs’ understanding of the world and explored how different persuasion strategies could modify LLMs’ worldviews. Different from our work, the aforementioned ToM and socialization benchmarks evaluate LLMs by placing them in a passive observer’s third-person perspective.
Necessity of developing LLMs’ ToM and socialization capabilities
Several studies (Gandhi et al., 2023; Kim et al., 2023) have shown that LLMs have poor reasoning performance and robustness on ToM and socialization tasks in a zero-shot setting, even with the current state-of-the-art GPT series (Achiam et al., 2023) model. With LLMs becoming increasingly integrated into our everyday lives, developing LLMs with ToM and socialization could be better at teaching us, learning from us, communicating with us, collaborating with us, and understanding us (Gandhi et al., 2021; 2023; Rabinowitz et al., 2018; Shu et al., 2021).
Ego-centric (First-person Perspective) Research
In the fields of computer vision and robotics, there has already been considerable research on a first-person perspective. For example, Cheng et al. (2023) explored whether vision-language models can ”Think from a First-person Perspective?” Huang et al. (2023) proposes the construction of embodied agents in a 3D world, which involves acquiring and processing first-person perspective images. Huang et al. (2024) built a bridge between third-person and first-person perspectives at the action level, while Dou et al. (2024) proposed a method designed to transform exocentric video-language data for egocentric video representation learning. However, research on first-person perspectives in the field of natural language processing remains unexplored.
Figure 2: The foundation, inspiration, and detailed methods for converting the third-person ToM benchmark into a first-person perspective.
LLMs for Interactive Scenarios
Some work focuses on designing interaction strategies to enable LLMs to gain more benefits during interactions. For example, Zhang et al. (2024a) proposed Agent-pro, Zhang et al. (2024b) introduced K-level reasoning, and Guo et al. (2023) put forward the Suspicion-Agent. Additionally, Li et al. (2023) explored Multi-LLM collaboration by informing LLMs of task rules through prompts. Park et al. (2023) introduced generative agents that can simulate human behavior. Bianchi et al. (2024) explored the social behavior of LLMs in negotiation scenarios. Fu et al. (2023) show LLMs can improve each other in a negotiation scenario. Fan et al. (2024) examined the capability of LLMs to make rational decisions in game theoretic scenarios.
3 EgoSocialArena
EgoSocialArena provides a systematic approach to convert third-person perspective ToM benchmarks into first-person perspective, constructs Rule-based Agents at different cognitive levels and Reinforcement Learning (RL) Agents, and includes various scenarios to evaluate LLMs’ ToM and socialization capabilities from a first-person perspective.
3.1 Converting Existing Third-person ToM Benchmarks to a First-person Perspective
Foundation and Inspiration
In LLM-based Agent applications, system message serves as a critical component, functioning to pre-set the model’s role and background. As illustrated in Figure 2(A), system message ”You are name and live in a town…” is used. Interestingly, in the domain of LLM self-awareness research (Laine et al., 2024), a similar linguistic construct is employed. As illustrated in Figure 2(B), researchers employ the pronoun ”you” to probe LLMs’ potential self-awareness. Inspired by and building upon studies in these two domains, we systematically modify system message, story, question, and answer options to transform third-person ToM benchmarks into a first-person perspective.
Conversion Method
As illustrated in Figure 2(C), unlike instructing LLMs in system message that ”you are a helpful assistant.”, we inform LLMs in system message that they have personally experienced certain social events, similar to deploy LLM-based Agent. As illustrated in Figure 2(D), we employ the pronoun ”you” to replace specific characters in stories and questions, thereby situating LLMs within particular roles. This approach enables the models to experience social events from a first-person perspective. The framing of questions is akin to that employed in self-awareness research. For modifications to answer options, consider LLMs answer from a first-person perspective, substituting ’I’ for specific character in the options. Higher-order ToM questions from third-person perspective benchmarks, after conversion, are still used to evaluate the ToM capabilities of LLMs from a first-person perspective. In contrast, first-order ToM questions, after conversion, form an assessment of LLMs’ mental states following social events, which we include in the scope of evaluating socialization capabilities.
3.2 Intriguing and Distinctive Social Situations
As illustrated in Figure 3, we design three particularly interesting scenarios—Counterfactual, New World, and Blackjack—all used to evaluate the socialization capabilities of LLMs.
Counterfactual
In real social situations, the rules of Rock-Paper-Scissors are: rock beats scissors, scissors beat paper, and paper beats rock. An LLM can relatively easily establish a belief based on this situation. In contrast, we define a counterfactual situation for the Rock-Paper-Scissors game (scissors beat rock, paper beats scissors, and rock beats paper) to explore whether an LLM can establish a belief that matches this counterfactual situation.
Figure 3: Evaluation examples for the Counterfactual, New World, and Blackjack scenarios in intriguing and distinctive social situations, as well as Number Guessing scenarios in the interactive environment.
New World
We design stories that describe new social world scenarios that are significantly different from the current social world. We aim to investigate whether LLMs can demonstrate cognitive flexibility and achieve substantive shifts in their understanding and reasoning about the social world.
Blackjack
In the single-turn card game scenario, it is necessary to analyze the game situation, i.e., analyze own cards in Self-belief and the opponent’s cards in External-belief. This forms a comprehensive mental estimation of the current game situation. We experiment with the blackjack card game, and its specific rules can be found in A.1.
3.3 Interactive Environment
In the interactive environment, many studies adopt two different LLMs to interact (e.g., GPT-3.5 vs GPT-4). However, we have observed a phenomenon in such experiments: ”Babysitting” weaker model distracts the stronger model. When comparing the ToM capabilities of the Deepseek-v2 model and the GPT-3.5 model from a first-person perspective, it is unreasonable to conduct experiments like GPT-3.5 vs GPT-4 and Deepseek-v2 vs GPT-4. Although both opponents are GPT-4, the babysitting phenomenon during the interaction process significantly interferes with the opponent’s output, leading to unfair comparisons. To ensure a fair and reasonable comparison, the opponent’s capability and behavioral strategies must be controlled to be stable during the interaction process. We propose two methods: constructing rule-based agents at different cognitive levels as opponents and training RL agents as opponents.
3.3.1 Rule-based Agents at Different Cognitive Levels
Agents’ actions at lower cognitive levels follow relatively simple and fixed rules. As the cognitive level increases, agents’ actions adhere to more complex rule patterns, exhibiting capabilities and behavior strategies that approximate human cognitive models. We establish rule-based agents at different cognitive levels as opponents and denote the action of LLM Agent and rule-based Agent as $a_{m}^{t}$ and $a_{o}^{t}$ in round $t$, respectively.
Scenario: Number Guessing
Level 1: $a_{o}^{t}=C$. In this pattern, we conduct experiments with the rule-based Agent’s actions remaining constant at 50. Level 2: $a_{o}^{t}=f(t)=50-5(t-1)$. In this pattern, we conduct experiments with the rule-based Agent’s action sequence of round 1: 50, round 2: 45, …, round 9: 10, round 10: 5, an arithmetic sequence with the first term 50 and a common difference of 5. Level 3: $a_{o}^{t}=f(a_{m}^{t-1},a_{o}^{t-1})=0.8\times\left(\frac{a_{m}^{t-1}+a_{o}^{t-1}}{2}\right)$. In this pattern, we conduct experiments with the rule-based Agent’s action copying the gold value from the previous round. The rules of number guessing can be found in A.1.
3.3.2 Reinforcement Learning Agents
In the Limit Texas Hold’em scenario, we train two reinforcement learning agents as opponents: Deep Q-network (DQN)-Aggressive (Mnih et al., 2015) and DQN-Conservative (Mnih et al., 2015). By adapting the reward function, RL agents are given different game personalities. For DQN-Aggressive, we encourage the action of raising and calling during the game. In contrast, for DQN-Conservative, we encourage the action of folding during the game. The rules of Limit Texas Hold’em can be found in A.1, and a specific example of the Limit Texas Hold’em scenario can be found in A.2.
3.3.3 Interaction
During the interaction between an LLM and a rule-based agent in Number Guessing scenario, as illustrated in Figure 3, each round will involve asking the LLM $\mathcal{M}$ ToM questions $\mathcal{Q}_{ng}$ regarding the opponent’s behavior:
$\mathcal{A}_{ng}^{t}=\mathcal{M}(\mathcal{Q}_{ng}^{t}\oplus\mathcal{H}^{\textless t}\oplus\mathcal{P}_{ng}),t\in\{1,2,\dots,N\}$ (1)
where $\mathcal{P}_{ng}$ is a prompt for leading to the answer, $t$ represents the current round number, <$\mathcal{H}^{\textless t}$ represents game history information, $\mathcal{A}_{ng}^{t}$ represents LLM’s response.
After multiple rounds of game interaction between the LLM and the RL Agent in Limit Texas Hold’em scenario, asking the LLM $\mathcal{M}$ ToM questions $\mathcal{Q}_{lth}$ regarding the opponent’s play style:
$\mathcal{A}_{lth}^{t}=\mathcal{M}(\mathcal{Q}_{lth}^{t}\oplus\mathcal{H}^{\textless t}\oplus\mathcal{P}_{lth}),t=N$ (2)
where $\mathcal{P}_{lth}$ is a prompt for leading to the answer, $t$ represents the current round number, <$\mathcal{H}^{\textless t}$ represents game history information, $\mathcal{A}_{lth}^{t}$ represents LLM’s response.
4 Data construction
4.1 Data Collection and Validation
The conversion of the third-person perspective ToM benchmark to the first-person perspective is achieved through GPT-4o, followed by manual verification and correction. The game hands for Limit Texas Hold’em and Blackjack card games are generated by RLcard (Zha et al., 2019). Additionally, we manually construct scenarios for both the new world and counterfactual situations. After the data collection, following Chen et al. (2024)’s method, we conduct two rounds of validation to ensure the data’s correctness and quality. In 1st round, author A would first complete all samples created by author B. For stories, questions, and answer options where there are disagreements, authors A and B would discuss and modify them to reach a consensus as much as possible. In 2nd round, for samples where consensus is still not reached, another author C would discuss with authors A and B to determine the final answer. After two rounds of discussion, the final average agreement reaches 97.6%.
4.2 Data Statistics
EgosocialArena includes two evaluation environments: static environment and interactive environment, with seven scenarios: Daily Life, Counterfactual, New World, Blackjack, Number-Guessing, and Limit Texas Hold’em, totaling 2,195 data entries. A comparison with existing ToM benchmarks is shown in Table 1.
ToM Benchmark Data Volume Evaluation Environment Number of Scenario Perspective
ToMi 999 Static 1 Third-person
HI-ToM 1200 Static 2 Third-person
FanToM 254 Static 2 Third-person
OpenToM 596 Static 3 Third-person
ToMBench 2860 Static 1 Third-person
EgosocialArena (Ours) 2195 Static and Dynamic 7 First-person
Table 1: Comparison of EgosocialArena with existing ToM benchmarks.
5 Experiments
5.1 Experimental Setup
We evaluate a total of 9 popular LLMs, including GPT-4o [^1] https://openai.com/index/hello-gpt-4o/, o1-preview [^2] https://openai.com/index/learning-to-reason-with-llms/, GPT-4-Turbo (Achiam et al., 2023), GPT-3.5-Turbo (Achiam et al., 2023), Claude-3.5-sonnet-20240620333https://www.anthropic.com/news/claude-3-5-sonnet, LLaMa-3-8B-Chat [^4] https://ai.meta.com/blog/meta-llama-3/, LLaMa-3-70B-Chat, LLaMa-3-8B-Instruct-Turbo, and LLaMa-3.1-405B-instruct-Turbo (Dubey et al., 2024). To account for the potential influence of model parameters and instruction tuning, we specifically compare LLaMa-3-8B-Chat with LLaMa-3-8B-Instruct-Turbo, as well as LLaMa-3-8B-Chat with LLaMa-3-70B-Chat.
To establish a human performance baseline, we recruit 10 graduate students, all of whom have received a good basic education and possess mature cognitive abilities, to complete responses to the questions in EgoSocialArena. The average accuracy of their responses will serve as the human performance baseline. No extra tutorials or examples are provided to ensure a fair comparison.
5.2 Evaluation Method
For Daily Life, New World, and CounterFactful scenarios, we present LLMs with a story, a question, and several options, then ask them to pick the correct answer. Using the accuracy of answering questions as the evaluation metric for these scenarios. In the interactive environments of Number Guessing and Texas Hold’em, the evaluation of these scenarios also has standard answers because we propose agents with stable capabilities and behavioral strategies as opponents of LLMs. For the Blackjack scenario, we conducted a manual evaluation. To ensure the quality of the manual evaluation, we measure the average consistency score between evaluators, which reached 96.3%.
Methods ToMI Number Guessing Texas Hold’em
Third-person First-person Level 1 Level 2 Level 3
Open-source Models
LLaMa-3-8B-Chat 50.6 66.2 0.0 0.0 0.0 48.0
LLaMa-3-70B-Chat 58.4 63.2 10.0 20.0 10.0 38.0
LLaMa-3-8B-Instruct 51.1 65.4 0.0 0.0 0.0 44.0
LLaMa-3.1-405B-Instruct 58.0 65.8 80.0 20.0 20.0 56.0
API-based Models
Claude-3-5-Sonnet 71.0 80.5 50.0 10.0 40.0 66.0
GPT-3.5-Turbo 45.5 51.9 10.0 10.0 0.0 56.0
GPT-4-Turbo 55.4 69.7 10.0 20.0 10.0 60.0
GPT-4o 64.1 71.0 10.0 40.0 10.0 62.0
o1-preview 71.9 77.5 90.0 90.0 90.0 72.0
Human
Human Performance 90.2 90.2 90.0 86.0 73.0 82.0
Table 2: The performance of ToM capabilities from the first-person perspective in open-source models and API-based models. The highest and second-highest scores among models and humans in each section are highlighted in blue and red, respectively.
5.3 Experimental Results
The performance of ToM and socialization capabilities from the LLMs’ first-person perspective is shown in Table 2 and 3, respectively. Based on the experimental results, we have obtained some key insights:
Performance Differences in LLMs’ Theory of Mind (ToM) Across Third-Person and First-Person Perspective
As shown in Table 2, all LLMs exhibited improved performance after the ToMI dataset was adapted from a third-person to a first-person perspective. Notably, the Claude and o1-preview models demonstrated significantly stronger ToM capabilities in the first-person perspective compared to other models. Except for GPT-3.5-Turbo, API-based models generally outperformed open-source models, including the recently released LLaMa-3.1-405B-Instruct. However, despite these improvements, there remains a substantial gap between the performance of all LLMs and that of human participants.
Methods New World Counterfact Blackjack ToMBench
All Desire Emotion Intention
Open-source Models
LLaMa-3-8B-Chat 6.7 71.0 84.5 67.4 71.5 73.0 50.0
LLaMa-3-70B-Chat 13.3 59.0 92.5 73.5 72.5 63.5 83.0
LLaMa-3-8B-Instruct 6.7 64.0 81.4 55.1 45.8 70.5 53.5
LLaMa-3.1-405B-Instruct 36.7 66.0 96.5 77.6 75.0 75.5 80.6
API-based Models
Claude-3-5-Sonnet 90.0 74.0 96.9 79.6 72.0 64.5 82.6
GPT-3.5-Turbo 13.3 37.0 89.7 72.5 70.0 86.0 68.5
GPT-4-Turbo 23.3 70.0 95.0 75.6 84.5 67.4 84.0
GPT-4o 36.7 52.0 96.5 85.7 78.0 91.0 96.5
o1-preview 86.7 90.0 97.2 84.8 73.0 91.0 98.0
Human
Human Performance 93.3 91.0 97.0 90.7 83.2 92.0 97.5
Table 3: The performance of socialization capabilities from the first-person perspective of open-source models and API-based models (after the ToMBench conversion, it also includes part of the ToM capability assessment, which we have consolidated here together) with the highest and second-highest scores among models and humans in each section are highlighted in blue and red, respectively.
The powerful capabilities of the o1-preview model are truly surprising
In the Number-Guessing scenario, almost all large language models perform poorly, even in the simplest Level 1 situation, which poses a significant challenge for humans as well. However, the recent o1-preview model has performed exceptionally well, demonstrating a cognitive level in this scenario that approaches or even surpasses that of humans.
The scaling up of open-source models or performing instruction fine-tuning has not yielded significant results
By comparing the performance of LLaMa-3-8B-Chat with LLaMa-3-70B-Chat, as well as LLaMa-3-8B-Chat with LLaMa-3-8B-Instruct models in Table 2, 3, it is observed that neither of these measures has notably improved the ToM and socialization capabilities of LLMs. To enhance LLMs’ ToM and socialization capabilities, innovative approaches will be needed in future research on LLMs.
The performance gap between open-source models and API-based models in socialization scenarios is significant
In counterfactual and New World scenarios, API-based models exhibit extremely flexible cognition, demonstrating a strong ability to adapt to social situations and showing performance comparable to humans. In contrast, open-source models perform poorly. For instance, the LLaMa-3-8B-Chat model only achieves an accuracy rate of 6.7% in New World scenarios, compared to 90.0% for the Claude model.
In the dimensions of desire and intention in social contexts, LLMs have shown potential
As illustrated in Table 3, although LLMs still lag behind humans in the overall dimension of mental states, their performance in the desire and intention dimensions is impressive. These findings can further inspire and serve as a foundation for the application of LLM-based agents in the real social world.
Figure 4: In the scenario of Number Guessing Level 2 (Arithmetic sequence), the belief state evolution pattern of GPT-4-Turbo regarding the opponent’s proposed numbers.
5.4 Qualitative Analysis
Analysis of Model Failure Causes and Behavioral Patterns in Number Guessing Scenario —— Mid-point Belief, Strange Guess and Get Back on Track
As shown in Figure 4, in the scenario of Number Guessing (Level 2: Arithmetic sequence), we thoroughly investigate the belief state evolution pattern of GPT-4-Turbo regarding the opponent’s proposed numbers. In round 1, with no available information, the GPT-4-Turbo model thinks the opponent will choose the number 50 within the range of 1-100. The same phenomenon is observed in the GPT-3.5-Turbo model, called ”mid-point belief”. Sometimes, the GPT-4-Turbo model continuously believes the opponent will choose progressively smaller numbers throughout the entire interaction, as depicted by the GPT-4-Turbo guess1 curve in Figure 4. Although this is very close to the gold number, it does not capture that the opponent’s chosen numbers form an arithmetic sequence. Another situation occurs when the GPT-4-Turbo model makes a ”strange guess” in the initial rounds, thinking the opponent will suddenly choose larger numbers. After several rounds, it captures that the opponent’s chosen numbers form an arithmetic sequence, called ”Get Back on Track”. Overall, despite the statistical results indicating that the GPT-4-Turbo model does not establish a belief regarding the Level 2 opponent in the Number Guessing scenario, the phenomena we observed suggest that it has started to grasp some patterns.
Analysis of our research and current studies centered on outcomes and benefits
Many current works focus on designing various strategies to improve the performance of LLMs in static or interactive environments. For example, making an LLM display anger to enhance its performance in negotiation scenarios, or using carefully crafted prompt engineering to analyze an opponent’s behavior from multiple perspectives to achieve higher payoffs. Unlike these approaches, we fundamentally explore the ToM capabilities of LLMs from a first-person perspective in various scenarios. Based on the behavior patterns of LLMs that we observed in our experiments, such as the excellent performance of the o1-preview and Claude models in the situated new world scenarios, it indicates that these models have significant potential for further exploration in fields like role-playing and simulation, as well as entering the real social world.
6 Conclusion
In this paper, considering the limitations of existing ToM and socialization benchmarks, the importance of first-person ToM and socialization capabilities in LLMs, and the natural approach of observing and understanding the world from an ego-centric first-person perspective for both humans and AI agents, we propose the EgoSocialArena framework. This framework is designed to comprehensively evaluate and probe the first-person ToM and socialization capabilities of LLMs in both static and interactive environments, covering multiple scenarios. To avoid the ”babysitting” problem during the evaluation process and achieve fair and comprehensive assessments, we construct rule-based agents at different cognitive levels and train RL agents. We collect a total of 2195 test data entries and test multiple advanced and popular LLMs, revealing some interesting and key insights, and highlighting a significant potential for enhancing the first-person ToM and socialization capabilities of LLMs, allowing them to truly enter the social world.
Appendix A Appendix
A.1 Game Rules
Blackjack
Blackjack, also known as 21, is a popular card game that involves a dealer and a player. Players must decide whether to hit or stand based on their own hand, the dealer’s face-up card, and the dealer’s one hidden card. The objective is to beat the dealer without exceeding 21 points.
Limit Texas Hold’em
The game commences with each player being dealt two private cards, which belong exclusively to the player and remain hidden from the others. Five community cards are then dealt face-up in a series of stages: a three-card Flop, followed by a single card on the Turn and another single card on the River. The player can choose from four actions: Fold, Check, Call, Raise. They aim to construct the best five-card poker hand possible using any combination of their private cards and community cards.
Number Guessing
Each player selects a number between 1 to 100. The objective is to select a number that is closest to 80% of the group’s average choice. The key idea is to guess how others will estimate the average and thus decide the number to submit.
A.2 Limit Texas Hold’em Scenario
Figure 5: Evaluation examples for the Limit Texas Hold’em Scenario.

### v2

@@
-Entering Real Social World! Benchmarking the Theory of Mind and Socialization Capabilities of LLMs from a First-person Perspective
+Entering Real Social World! Benchmarking the Social Intelligence of Large Language Models from a First-person Perspective
 Abstract
-In the social world, humans possess the capability to infer and reason about others’ mental states (such as emotions, beliefs, and intentions), known as the Theory of Mind (ToM).
-Simultaneously, humans’ own mental states evolve in response to social situations, a capability we refer to as socialization.
-Together, these capabilities form the foundation of human social interaction.
-In the era of artificial intelligence (AI), especially with the development of large language models (LLMs), we raise an intriguing question: How do LLMs perform in terms of ToM and socialization capabilities?
-And more broadly, can these AI models truly enter and navigate the real social world?
-Existing research evaluating LLMs’ ToM and socialization capabilities by positioning LLMs as passive observers from a third-person perspective, rather than as active participants.
-However, compared to the third-person perspective, observing and understanding the world from an ego-centric first-person perspective is a natural approach for both humans and AI agents.
-The ToM and socialization capabilities of LLMs from a first-person perspective, a crucial attribute for advancing embodied AI agents, remain unexplored.
-To answer the aforementioned questions and bridge the research gap, we introduce EgoSocialArena, a novel framework designed to evaluate and investigate the ToM and socialization capabilities of LLMs from a first-person perspective.
-It encompasses two evaluation environments: static environment and interactive environment, with seven scenarios: Daily Life, Counterfactual, New World, Blackjack, Number Guessing, and Limit Texas Hold’em, totaling 2,195 data entries.
-With EgoSocialArena, we have conducted a comprehensive evaluation of nine advanced LLMs and observed some key insights regarding the future development of LLMs as well as the capabilities levels of the most advanced LLMs currently available.
+Social intelligence is built upon three foundational pillars: cognitive intelligence, situational intelligence, and behavioral intelligence.
+As large language models (LLMs) become increasingly integrated into our social lives, understanding, evaluating, and developing their social intelligence are becoming increasingly important.
+While multiple existing works have investigated the social intelligence of LLMs, (1) most focus on a specific aspect, and the social intelligence of LLMs has yet to be systematically organized and studied; (2) position LLMs as passive observers from a third-person perspective, such as in Theory of Mind (ToM) tests.
+Compared to the third-person perspective, ego-centric first-person perspective evaluation can align well with actual LLM-based Agent use scenarios.
+(3) a lack of comprehensive evaluation of behavioral intelligence, with specific emphasis on incorporating critical human-machine interaction scenarios.
+In light of this, we present EgoSocialArena, a novel framework grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence, aimed to systematically evaluate the social intelligence of LLMs from a first-person perspective.
+With EgoSocialArena, we conduct a comprehensive evaluation of eight prominent foundation models, even the most advanced LLMs like o1-preview lag behind human performance by 11.0 points.
+Our code and data are available at https://github.com/gyhou123/EgoSocialArena.
 1 Introduction
-In the complex social interactions of humans, two fundamental cognitive capabilities play crucial roles: Theory of Mind (ToM) and Socialization.
-ToM is a fundamental psychological process, defined as the capacity to reason about others’ mental states – beliefs, intents, desires, emotions, knowledge, etc (Premack & Woodruff, 1978; Ma et al., 2023; Chen et al., 2024).
-Socialization refers to the capability for own mental state evolution in response to social situations.
-As illustrated in Figure 1(A), When the little boy receives a birthday gift or achieves good grades, he feels very happy.
-These intertwined capabilities are central to humans’ social life.
-With the advent of the era of LLMs, powerful models like GPT-4 (Achiam et al., 2023) and Claude (Anthropic, 2024) have demonstrated remarkable competence in multiple tasks and domains.
-It raises intriguing questions: How do LLMs perform in terms of ToM and socialization capabilities?
-And more broadly, can these AI models truly enter and navigate the real social world to achieve efficient human-computer interaction?
-To evaluate the ToM and socialization capabilities of LLMs, multiple benchmarks have been proposed, such as SocialIQA (Sap et al., 2022), NormBank (Ziems et al., 2023), BigToM (Gandhi et al., 2023), FanToM (Kim et al., 2023), HI-ToM (Wu et al., 2023), OpenToM (Xu et al., 2024), NegotiationToM (Chan et al., 2024), and ToMBench (Chen et al., 2024).
-However, as illustrated in Figure 1(B), these benchmarks evaluate LLMs’ ToM and socialization capabilities by positioning LLMs as passive observers from a third-person perspective rather than as active participants.
-We propose two key points: (1) Observing and understanding the world from an ego-centric first-person perspective is a natural approach for both humans and AI agents.
-(2) Evaluating LLM’s ToM and socialization capabilities from an ego-centric first-person perspective measures whether LLMs can enter the real social world to achieve efficient human-computer interaction, providing indispensable resources for future research in the field of Embodied Artificial Intelligence (EAI).
-Figure 1: (A): ToM (others mental states), Socialization (own mental states).
-(B): LLM acts as a passive observer to analyze the mental states of characters within a story from a third-person perspective.
-(C): Evaluation Scenarios in EgosocialArena.
-(D):Rule-based Agents and RL Agents within Interactive environment.
-In this paper, we present EgosocialArena, a novel framework for evaluating and investigating LLMs’ ToM and socialization capabilities from a first-person perspective.
-We propose a structured taxonomy to guide the development of EgoSocialArena, encompassing three key components: (1) We propose a systematic methodology to transform two existing third-person ToM benchmarks, ToMI and ToMbench, into a first-person perspective.
-We present a comprehensive argument to justify the validity and rationale behind this transformation.
-The higher-order ToM questions in a third-person perspective are transformed to evaluate the first-person ToM capabilities of LLMs.
-In contrast, first-order ToM questions in a third-person perspective are transformed to evaluate LLMs’ mental states in response to social events, i.e., socialization capabilities.
-Detailed descriptions can be found in 3.1.
-(2) We investigate the socialization capabilities of LLMs in intriguing and distinctive social situations, including counterfactual, new world, and blackjack game scenarios, as illustrated in Figure 1(C).
-(3) We investigate the ToM capabilities of LLMs within interactive environments, including belief about the opponent’s behavior patterns in number-guessing scenarios and modeling of opponent’s play styles in Texas Hold’em poker game scenarios.
-In interactions between models of varying strengths, such as GPT-3.5 and GPT-4, we observe a ”babysitting” issue, where the weaker model can negatively impact the stronger one, diminishing its performance.
-To mitigate this issue and enable fair evaluation, as illustrated in Figure 1(D), we construct rule-based agents at different cognitive levels and train Reinforcement Learning (RL) agents to provide stable capabilities and behavioral strategies.
-Overall, EgoSocialArena encompasses two evaluation environments: static environment and interactive environment, with seven scenarios: Daily Life, Counterfactual, New World, Blackjack, Number Guessing, and Limit Texas Hold’em, totaling 2,195 data entries.
-We conduct extensive experiments on EgoSocialArena to evaluate 9 foundational models known for their leading performance across multiple tasks and domains.
-This set includes five API-based models (i.e., o1-preview, GPT-4o, GPT-4-Turbo, GPT-3.5-Turbo, and claude-3-5-sonnet-20240620) and four open-source models (LLaMa-3-8B-Chat, LLaMa-3-70B-Chat, LLaMa-3-8B-Instruct, and LLaMa-3.1-405B-Instruct).
-We establish a human performance baseline by engaging qualified human annotators.
-Our experimental results reveal several interesting and critical insights:
-  1. 1.
-    Although LLMs currently perform behind humans in most scenarios, they have shown significant potential in a few specific cases.
-  2. 2.
-    The ToM capabilities of LLMs show significant differences between a third-person perspective and a first-person perspective.
-  3. 3.
-    The powerful capabilities of the o1-preview model are truly surprising.
-  4. 4.
-    The scaling up of open-source models or performing instruction fine-tuning has not yielded significant results in terms of ToM and socialization capabilities.
-  5. 5.
-    The performance gap between open-source models and API-based models in socialization scenarios is significant.
-2 Related Work
-Existing ToM Benchmarks
-Previous evaluations for the ToM of LLMs primarily focus on testing models using narrative stories, also referred to as reading comprehension scenarios.
-Specifically, Le et al.
-(2019) proposed the ToMi benchmark based on the classic Sally-Anne test.
-Wu et al.
-(2023) introduced the HI-ToM benchmark, which focuses on higher-order belief reasoning and sets up scenarios where agents can communicate with each other.
-Gandhi et al.
-(2023) proposed BigToM, which presents a framework for designing a ToM benchmark from synthetic templates for evaluating different aspects of LLMs’ ToM capabilities (e.g., desire and belief).
-Xu et al.
-(2024) introduced OpenToM, which assigns personalities to agents in the stories and ensures that the storylines are more reasonable and logical.
-Chen et al.
-(2024) proposed ToMBench, which systematically evaluates LLMs across all dimensions of ToM capabilities.
-Unlike the above methods that require LLMs to read stories and answer related questions, some studies evaluate LLMs’ performance by inputting dialogues to them.
-Kim et al.
-(2023) proposed FanToM, which tests LLMs on their ability to infer the mental states of characters in everyday conversations.
-Chan et al.
-(2024) introduced NegotiationToM, which restricts the dialogue content to negotiation scenarios.
-Existing Socialization Benchmarks
+Social intelligence, i.e., the ability to understand and reason about the mental states of others (cognitive intelligence), awareness and adaptation to the social context (situational intelligence), and effective interaction with others (behavioral intelligence), is a form of advanced intelligence that naturally develops during human growth Thorndike (1921); Hunt (1928); Premack and Woodruff (1978); Hou et al.
+(2024); Li et al.
+(2024).
+Imagine the future where robots powered by large language models (LLMs) enter our social world, communicating with us empathetically, supporting us in living better, and making great contributions to society.
+This is a wonderful vision and highlights the importance and significance of understanding, evaluating, and developing the social intelligence of LLMs.
+Numerous datasets have been curated to assess the social intelligence of LLMs, such as ToMI Le et al.
+(2019), BigToM Gandhi et al.
+(2023), FanToM Fan et al.
+(2024), HI-ToM Wu et al.
+(2023), OpenToM Xu et al.
+(2024), and ToMBench Chen et al.
+(2024b) for evaluating Theory of Mind (ToM) capabilities of LLMs, focusing on reasoning about the mental states of others; SocialIQa Sap et al.
+(2022) and NormBank Ziems et al.
+(2023) for evaluating LLMs’ understanding of social contexts; SOTOPIA Zhou et al.
+(2023) and LLMArena Chen et al.
+(2024a) for evaluating LLMs’ behavior and interaction capabilities in social goal-driven and gaming scenarios.
+However, as illustrated in Figure 1(A), these existing works each focus on a specific aspect of social intelligence, such as ToM tests corresponding to cognitive intelligence, and the social intelligence of LLMs has not yet been systematically organized and studied.
+On the other hand, as illustrated in Figure 1(B), these existing works evaluate LLMs’ ToM and social context understanding abilities by positioning LLMs as passive observers from a third-person perspective.
+We propose two key points: (1) The third-person perspective involves making LLMs engage in "armchair theorizing" that isn’t aligned with real LLM-based Agent use scenarios.
+This kind of evaluation isn’t accurate enough.
+(2) Ego-centric first-person perspective evaluation can align well with actual LLM-based Agent use scenarios, allowing us to better and more thoroughly understand their performance in human society.
+Figure 1: (A): Datasets related to social intelligence over time in the Era of LLMs (a non-exhaustive visualization due to space constraints).
+(B): LLM acts as a passive observer to analyze mental states of characters within a story from a third-person perspective.
+(C): Main direction of existing work on the behavioral intelligence of LLMs.
+Moreover, as illustrated in Figure 1(C), when evaluating the behavioral and interactive capabilities of LLMs, existing work like LLMArena propose various game environments and have different LLMs interact to see who wins and loses.
+Compared to having two LLMs play games to determine winners and losers, exploring LLM’s performance in human-machine interaction is more meaningful.
+Additionally, many works, such as Hypothetical Minds Cross et al.
+(2024) and SOTOPIA-Pi Wang et al.
+(2024), focus on proposing various strategies, such as prompt-based methods or behavior cloning, to enhance the performance of LLMs in interactive environments like Melting 2.0 Agapiou et al.
+(2022) and SOTOPIA.
+However, there is still a lack of comprehensive evaluation of behavioral intelligence for current mainstream LLMs.
+In this paper, we present EgoSocialArena, a novel framework designed to systematically evaluate the social intelligence of LLMs from a first-person perspective.
+The development of EgoSocialArena is grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence: (1) Cognitive Intelligence: we propose a complete and generalizable workflow to transform existing static third-person ToM benchmarks into a first-person perspective.
+Additionally, we have newly developed a dynamic cognitive assessment in multi-turn interactive scenarios.
+(2) Situational Intelligence: Imagine an LLM-based Agent entering our social world - how would it respond emotionally when receiving praise or gifts [^1] This might be related to self-awareness, but the focus could be shifted more towards the application situations.?
+We have newly developed an assessment for such real-world social situations.
+Additionally, we have also developed assessments for counterfactual situations and parallel world situations.
+(3) Behavioral Intelligence: we incorporate existing cooperative and adversarial game environments, as well as social goal-driven interactive dialogue environments, to comprehensively evaluate the behavioral intelligence of LLMs.
+Overall, as illustrated in Figure 2, EgoSocialArena encompasses the evaluation of cognitive, situational, and behavioral intelligence, with eight scenarios: static cognition, dynamic cognition evolution, real-world social situation, counterfactual situation, parallel world situation, cooperative game, adversarial game, and social goal-driven human-machine interactive dialogue environment, comprising a total of 2245 data entries.
+Figure 2: Examples of eight scenarios in EgoSocialArena.
+We conduct extensive experiments on EgoSocialArena to evaluate 8 foundational models known for their leading performance across multiple tasks and domains.
+This set includes five API-based models (i.e., o1-preview, GPT-4o, GPT-4-Turbo, GPT-3.5-Turbo, and claude-3-5-sonnet-20240620) and three open-source models (LLaMa-3-8B-Chat, LLaMa-3-70B-Chat, and LLaMa-3.1-405B-Instruct).
+We establish a human performance baseline by engaging qualified human annotators with a college degree or higher.
+Our experimental results reveal several interesting and critical insights: (1) The o1-preview model achieved the highest score of 80.6 among all models, surpassing human performance in dynamic cognition and adversarial game scenarios.
+Nevertheless, an 11.0 gap in overall accuracy remains when compared to the human baseline, leaving plenty of room for model improvement.
+Our in-depth analysis reveals that the superiority of o1-preview is mainly attributed to its powerful logical reasoning and mathematical abilities (keenly uncovering deeper patterns behind the data).
+(2) Comparing the performance of LLaMA3-8B-Chat and LLaMA3-70B-Chat models shows that simply scaling model size does not significantly help improve the social intelligence of LLMs.
+(3) Compared to the third-person perspective, LLMs show significantly improved ToM reasoning ability when operating from a first-person perspective.
+2 EgoSocialArena
+2.1 Cognitive Intelligence
+In the static cognition scenario, we convert the existing third-person ToMI benchmark to a first-person perspective.
+In the dynamic cognition evolution scenario, we construct opponents with various behavioral strategies, including rule-based agents at different cognitive levels and Reinforcement Learning (RL) agents, to explore how LLMs can form beliefs about opponents’ behavioral strategies during multi-round interactions.
+2.1.1 Static Cognition — Converting Existing Third-person ToM Benchmarks to a First-person Perspective
+Foundation and Inspiration
+In LLM-based Agent applications, system message serves as a critical component, functioning to pre-set the model’s role and background.
+As illustrated in Figure 3(A), system message "You are name and live in a town…" is used.
+Interestingly, in the domain of LLM self-awareness research (Laine et al., 2024), a similar linguistic construct is employed.
+As illustrated in Figure 3(B), researchers employ the pronoun "you" to probe LLMs’ potential self-awareness.
+Inspired by and building upon studies in these two domains, we systematically modify system message, story, question, and answer options to transform third-person ToM benchmarks into a first-person perspective.
+Figure 3: The foundation, inspiration, and detailed methods for converting the third-person ToM benchmark into a first-person perspective.
+Conversion Method
+As illustrated in Figure 3(C), unlike instructing LLMs in system message that "you are a helpful assistant.", we inform LLMs in system message that they have personally experienced certain social events, similar to deploy LLM-based Agent.
+As illustrated in Figure 3(D), we employ the pronoun "you" to replace specific characters in stories and questions, thereby situating LLMs within particular roles.
+This approach enables the models to experience social events from a first-person perspective.
+The framing of questions is akin to that employed in self-awareness research.
+2.1.2 Dynamic Cognition Evolution — Number Guessing (G0.8A)
+Scenario: G0.8A
+Each player selects a number between 1 to 100. The objective is to select a number that is closest to 80% of the group’s average choice.
+Rule-based Agents at Different Cognitive Levels
+Agents’ actions at lower cognitive levels follow relatively simple and fixed rules.
+As the cognitive level increases, agents’ actions adhere to more complex rule patterns, exhibiting capabilities and behavior strategies that approximate human cognitive models.
+We establish rule-based agents at different cognitive levels as opponents and denote the action of LLM Agent and rule-based Agent as $a_{m}^{t}$ and $a_{o}^{t}$ in round $t$, respectively.
+Level 1: $a_{o}^{t}=C$.
+In this pattern, we conduct experiments with the rule-based Agent’s actions remaining constant at 50.
+Level 2: $a_{o}^{t}=f(t)=50-5(t-1)$.
+In this pattern, we conduct experiments with the rule-based Agent’s action sequence of round 1: 50, round 2: 45, …, round 9: 10, round 10: 5, an arithmetic sequence with the first term 50 and a common difference of 5.
+Level 3: $a_{o}^{t}=f(a_{m}^{t-1},a_{o}^{t-1})=0.8\times\left(\frac{a_{m}^{t-1}+a_{o}^{t-1}}{2}\right)$.
+In this pattern, we conduct experiments with the rule-based Agent’s action copying the gold value from the previous round.
+2.1.3 Dynamic Cognition Evolution — Limit Texas Hold’em
+Scenario: Limit Texas Hold’em
+The game commences with each player being dealt two private cards Five community cards are then dealt face-up in a series of stages: a three-card Flop, followed by a single card on the Turn and another single card on the River.
+The player can choose from four actions: Fold, Check, Call, Raise.
+Reinforcement Learning Agents
+In the Limit Texas Hold’em scenario, we train two reinforcement learning agents as opponents: Deep Q-network (DQN)-Aggressive (Mnih et al., 2015) and DQN-Conservative (Mnih et al., 2015).
+By adapting the reward function, RL agents are given different game personalities.
+For DQN-Aggressive, we encourage the action of raising and calling during the game.
+In contrast, for DQN-Conservative, we encourage the action of folding during the game.
+A specific example of the Limit Texas Hold’em scenario can be found in Appendix B.
+2.2 Situational Intelligence
+2.2.1 Real-World Social Situation
+By filtering data from SocialIQa and ToMBench and using the transformation method mentioned in section 2.1.1, we evaluate the mental states of LLMs’ self after experiencing certain social events from a first-person perspective.
+2.2.2 Counterfactual Situation
+The conventional rules of Rock-Paper-Scissors (RPS) are: rock beats scissors, scissors beat paper, and paper beats rock.
+An LLM can relatively easily adapt to this situation.
+In contrast, we define a counterfactual situation for the RPS game (scissors beat rock, paper beats scissors, and rock beats paper) to explore whether an LLM can achieve situational adaptation.
+In addition to constructing counterfactual situations like RPS games, we also construct counterfactual situations based on physical facts, chemical facts, biological facts, traffic rules, social etiquette knowledge, etc.
+2.2.3 Parallel World Situation
+We design narratives depicting parallel social world that differ significantly from our current social world.
+We aim to investigate whether LLMs can demonstrate situational adaptation to these alternative worlds.
+2.3 Behavioral Intelligence
+2.3.1 Adversarial Game
+Blackjack, also known as 21 points, is a card game that involves a dealer and a player.
+The player must decide whether to hit or stand based on own hand, the dealer’s face-up card, and the dealer’s one hidden card.
+The objective is to beat the dealer without exceeding 21 points.
+We evaluate the win rate of LLMs as a player in this scenario.
+2.3.2 Cooperative game
+Defuse Bomb: Three LLMs emulate specialists in a team to defuse bombs.
+Bombs are distributed across $n$ rooms, whether the rooms are interconnected can be set manually.
+Each bomb exhibits unique phase sequences in $m$ colors, requiring the correct order of wire cutters for defusing.
+Team members start with different colored cutters and must coordinate and synchronize efforts for efficiency.
+We create 5 different map environments, each containing 5 bombs.
+Following Li et al.
+(2023), each successfully defused bomb awards the team 10 points per processed phase.
+We measure collaboration efficiency by calculating the score a team composed of three LLMs can achieve within 10 rounds.
+2.3.3 Social-goal Driven Human-Machine Interactive Dialogue
+With an open-ended social interaction environment SOTOPIA Zhou et al.
+(2023), which assigns a social goal and character to each agent involved.
+We focus on a comprehensive evaluation of interactions between current mainstream LLMs and humans, aiming to provide a more intuitive comparison of behavioral differences between humans and LLMs in social goal-driven interactive dialogue.
+We use the goal completion metric to quantitatively express this difference.
+3 Data Collection, Validation and Statistics
+The conversion of the third-person perspective to the first-person perspective is achieved through GPT-4o, followed by manual verification and correction.
+The game hands for Limit Texas Hold’em and Blackjack card games are generated by RLcard (Zha et al., 2019).
+Defuse bomb environment is based on gym API (Brockman, 2016) and a text interface.
+Additionally, we manually construct datasets for both the parallel world and counterfactual situations.
+After the data collection, following Chen et al.
+(2024b)’s method, we conduct two rounds of validation to ensure the data’s correctness and quality.
+In 1st round, author A would first complete all samples created by author B.
+For stories, questions, and answer options where there are disagreements, authors A and B would discuss and modify them to reach a consensus as much as possible.
+In 2nd round, for samples where consensus is still not reached, another author C would discuss with authors A and B to determine the final answer.
+After two rounds of discussion, the final average agreement reaches 97.6%.
+Data statistics of EgoSocialArena are shown in Table 1.
+Statistics Number Data Source
+Cognitive Intelligence 1235
+-Static Cognition 1155 Convertion
+-Dynamic Cogntion Evolution-N0.8A 30 Newly Created
+-Dynamic Cognition Evolution-Texas 50 Newly Created
+Situational Intelligence 675
+-Parallel World Situation 90 Newly Created
+-Counterfactual Situation 100 Newly Created
+-Real Social World Situation 485 Filter, Convertion
+Behavioral Intelligence 335
+-Adversarial Game 300 Existing
+-Cooperative Game 15 Existing
+-Social Goal 20 Existing
+Table 1: Data Statistics of EgoSocialArena.
+4 Experiments
+4.1 Experimental Setup
+We evaluate a total of eight prominent foundation LLMs, including GPT-4o [^2] https://openai.com/index/hello-gpt-4o/, o1-preview [^3] https://openai.com/index/learning-to-reason-with-llms/, GPT-4-Turbo (Achiam et al., 2023), GPT-3.5-Turbo (Achiam et al., 2023), Claude-3.5-sonnet-20240620444https://www.anthropic.com/news/claude-3-5-sonnet, LLaMa-3-8B-Chat [^5] https://ai.meta.com/blog/meta-llama-3/, LLaMa-3-70B-Chat, and LLaMa-3.1-405B-instruct-Turbo (Dubey et al., 2024).
+To account for the potential influence of model parameters, we specifically compare LLaMa-3-8B-Chat with LLaMa-3-70B-Chat.
+To establish a human performance baseline, we recruit 10 graduate students, all of whom have received a good basic education and possess mature cognitive abilities, to complete responses to the questions in EgoSocialArena.
+The average accuracy of their responses will serve as the human performance baseline.
+No extra tutorials or examples are provided to ensure a fair comparison.
+In the behavioral intelligence scenario, we similarly have these students participate in Adversarial Games and Cooperative Games, recording their average performance.
+For Social-Goal Driven Dialogue scenario, we use the performance of human interactions with GPT-4o as the baseline, given that GPT-4o is the best-performing LLM for this task.
+4.2 Evaluation Method
+For the evaluation of static cognition and situational intelligence, we present LLMs with a story, a question, and several options, then ask them to pick the correct answer.
+Using the accuracy of answering questions as the evaluation metric for these scenarios.
+For the evaluation of dynamic cognition evolution, these scenarios also has standard answers.
+For the adversarial and cooperative game scenario, we consider the win rate and team scores.
+For the Social-goal driven interactive dialogue, we use GPT-4 to automatically evaluate the performance of humans and LLMs in terms of goal completion during their interactions.
+Methods Cognitive Intelligence Static Cognition Dynamic Cognition-G0.8A Dynamic Cogntion Third-person First-person $\Delta$ Level 1 Level 2 Level 3 Limit Texas Open-source Models LLaMa-3-8B-Chat 50.6 66.2 +15.6 0.0 0.0 0.0 48.0 LLaMa-3-70B-Chat 58.4 63.2 +4.8 10.0 20.0 10.0 38.0 LLaMa-3.1-405B-Instruct 58.0 65.8 +7.8 80.0 20.0 20.0 56.0 API-based Models Claude-3-5-Sonnet 71.0 80.5 +9.5 50.0 10.0 40.0 66.0 GPT-3.5-Turbo 45.5 51.9 +6.4 10.0 10.0 0.0 56.0 GPT-4-Turbo 55.4 69.7 +14.3 10.0 20.0 10.0 60.0 GPT-4o 64.1 71.0 +6.9 10.0 40.0 10.0 62.0 o1-preview 71.9 77.5 +5.6 90.0 90.0 90.0 72.0 Human Human Performance 90.2 90.2 0.0 90.0 86.0 73.0 82.0
+Methods Situational Intelligence Behavioral Intelligence AVG Parallel World Counterfact Real-World Adversarial Cooperative Social Goal Open-source Models LLaMa-3-8B-Chat 6.7 71.0 67.2 51.3 49.7 22.5 34.8 LLaMa-3-70B-Chat 13.3 59.0 73.2 45.0 53.3 25.5 37.3 LLaMa-3.1-405B-Instruct 36.7 66.0 77.3 52.3 65.2 34.0 52.1 API-based Models Claude-3-5-Sonnet 90.0 74.0 79.8 55.0 94.8 50.5 62.8 GPT-3.5-Turbo 13.3 37.0 72.2 46.7 50.3 33.0 34.6 GPT-4-Turbo 23.3 70.0 75.7 54.7 75.6 52.0 47.4 GPT-4o 36.7 52.0 85.8 54.0 80.8 53.0 50.5 o1-preview 86.7 90.0 84.7 56.7 96.3 52.5 80.6 Human Human Performance 93.3 91.0 90.7 56.3 100.0 64.5 91.6
+Table 2: Performance of cognitive, situational, and behavioral intelligence from first-person perspective of eight LLMs.
+Highest and second-highest scores among LLMs and humans in each scenario are highlighted in red and blue, respectively.
+AVG represents the average value of cognitive, situational, and behavioral intelligence performance.
+4.3 Main Results
+As shown in Table 2, the o1-preview model achieved the highest score of 80.6 among all models, surpassing human performance in dynamic cognition and adversarial game scenarios.
+Nevertheless, an 11.0 gap in overall performance remains when compared to the human baseline, leaving plenty of room for model improvement.
+The second-best performer is the claude-3-5-sonnet model, which demonstrate impressive results in the static cognition and parallel world scenarios.
+The GPT-4o model performed well in the Real Social World Situation and Social Goal-Driven interactive dialogue scenarios, likely due to being trained with a substantial amount of human feedback.
+Overall, the performance of open-source models lags significantly behind that of API-based models and most models still exhibit a large performance gap compared to humans.
+For instance, the LLaMa-3-8B-Chat model achieved an overall score of 34.8, significantly lower than the human performance of 91.6.
+4.4 In-Depth Analysis
+Performance Differences in LLMs’ ToM Capabilities Across Third-Person and First-Person Perspective
+As shown in Table 2, all LLMs exhibited improved performance after the ToMI benchmark is converted from a third-person to a first-person perspective.
+The Llama3-8B-Chat model achieved the largest improvement of +15.6.
+Notably, the claude and o1-preview models demonstrated significantly stronger ToM capabilities in the first-person perspective compared to other models.
+Except for GPT-3.5-Turbo, API-based models generally outperformed open-source models, including the recently released LLaMa-3.1-405B-Instruct.
+However, despite these improvements, there remains a substantial gap between the performance of all LLMs and human baselines.
+The scaling up of open-source models has not yielded significant results
+By comparing the performance of LLaMa-3-8B-Chat with LLaMa-3-70B-Chat in Table 2, we observe that although the model size increased significantly, the overall performance on social intelligence improved by only +2.5.
+We further explore the scaling effects of increasing the size of the LLaMa-3 model on GSM8K (Cobbe et al., 2021) and MMLU (Chung et al., 2024) tasks, finding improvements of +12.9 and +13.4, respectively, as illustrated in Figure 4.
+Figure 4: Left: performance evolution corresponding to scaling up LLaMa-3 model size across different task domains. Right: o1-preview model’s output in dynamic cognition evolution—G0.8A scenario.
+The powerful mathematical capabilities of the o1-preview model are truly surprising
+In the dynamic cognition evolution-G0.8A scenario, almost all LLMs perform poorly, even in the simplest level 1 situation, which poses a significant challenge for humans as well.
+However, the recent o1-preview model has performed exceptionally well, we analyze its outputs and find that it is highly sensitive to numbers and can capture the correlations between numbers and the underlying patterns behind them, as illustrated in Figure 4.
+Therefore, when humans are unable to perceive these numerical patterns, the o1-preview model, based on its powerful mathematical capabilities, perceives things that humans have not detected.
+Mid-point Belief, Strange Guess and Get Back on Track
+As shown in Figure 5, in the scenario of dynamic cognition G0.8A Level 2 (Arithmetic sequence), we thoroughly investigate the belief state evolution pattern of GPT-4-Turbo regarding the opponent’s proposed numbers.
+In round 1, with no available information, the GPT-4-Turbo model thinks the opponent will choose the number 50 within the range of 1-100.
+The same phenomenon is observed in the GPT-3.5-Turbo model, called "mid-point belief".
+Sometimes, the GPT-4-Turbo model continuously believes the opponent will choose progressively smaller numbers throughout the entire interaction, as depicted by the GPT-4-Turbo guess1 curve in Figure 5.
+Although this is very close to the gold number, it does not capture that the opponent’s chosen numbers form an arithmetic sequence.
+Another situation occurs when the GPT-4-Turbo model makes a "strange guess" in the initial rounds, thinking the opponent will suddenly choose larger numbers.
+After several rounds, it captures that the opponent’s chosen numbers form an arithmetic sequence, called Get Back on Track.
+Overall, despite the statistical results indicating that the GPT-4-Turbo model does not establish a belief regarding the Level 2 opponent in the G0.8A scenario, the phenomena we observed suggest that it has started to grasp some patterns.
+The belief information for all models across all rounds can be found in Appendix C.
+Figure 5: In the scenario of G0.8A Level 2 (Arithmetic sequence), the belief state evolution pattern of GPT-4-Turbo regarding the opponent’s proposed numbers.
+5 Conclusion
+In this paper, considering the social intelligence of LLMs has yet to be systematically organized and studied, ego-centric first-person perspective evaluation can align well with actual LLM-based Agent use scenarios, incorporating human-machine interaction scenario evaluation is critical and the natural approach of observing and understanding the world from an ego-centric first-person perspective for both humans and LLM-based agents, we propose the EgoToMArena framework.
+This framework is grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence, with eight scenarios: static, dynamic cognition; real-world, counterfactual, parallel world situation; cooperative, adversarial game, and social goal-driven human-machine interactive environment, aimed to systematically evaluate the social intelligence of LLMs from a first-person perspective.
+We conduct extensive experiments and observe some key insights regarding the future development of LLMs as well as the capabilities levels of the most advanced LLMs currently available.
+Limitations
+There are three major limitations in our study.
+(1) Our study only involves the text modality and does not utilize ego-centric images and videos.
+The social intelligence of Vision-Language Models from a first-person perspective is very important, and we will leave this for future research.
+(2) Due to the constraint of computing resources and budget, we only evaluate eight prominent foundation LLMs, While we believe that the selected LLMs are representative.
+(3) Our study evaluates the social intelligence of LLMs from a first-person perspective, a deeper interpretation of these evaluation results from the perspective of explainability research would be more beneficial for the development of LLMs’ social intelligence.
+Appendix
+Appendix A Related Works
+Ego-centric (First-person Perspective) Research
+In the fields of computer vision and robotics, there has already been considerable research on a first-person perspective.
+For example, Cheng et al.
+(2023) explored whether vision-language models can "Think from a First-person Perspective?" Huang et al.
+(2023) proposes the construction of embodied agents in a 3D world, which involves acquiring and processing first-person perspective images.
+Huang et al.
+(2024) built a bridge between third-person and first-person perspectives at the action level, while Dou et al.
+(2024) proposed a method designed to transform exocentric video-language data for egocentric video representation learning.
+However, research on first-person perspectives in the field of natural language processing remains unexplored.
+Datasets Related to Social Intelligence
 Sap et al.
@@
 (2023) studied LLMs’ understanding of the world and explored how different persuasion strategies could modify LLMs’ worldviews.
-Different from our work, the aforementioned ToM and socialization benchmarks evaluate LLMs by placing them in a passive observer’s third-person perspective.
-Necessity of developing LLMs’ ToM and socialization capabilities
-Several studies (Gandhi et al., 2023; Kim et al., 2023) have shown that LLMs have poor reasoning performance and robustness on ToM and socialization tasks in a zero-shot setting, even with the current state-of-the-art GPT series (Achiam et al., 2023) model.
-With LLMs becoming increasingly integrated into our everyday lives, developing LLMs with ToM and socialization could be better at teaching us, learning from us, communicating with us, collaborating with us, and understanding us (Gandhi et al., 2021; 2023; Rabinowitz et al., 2018; Shu et al., 2021).
-Ego-centric (First-person Perspective) Research
-In the fields of computer vision and robotics, there has already been considerable research on a first-person perspective.
-For example, Cheng et al.
-(2023) explored whether vision-language models can ”Think from a First-person Perspective?” Huang et al.
-(2023) proposes the construction of embodied agents in a 3D world, which involves acquiring and processing first-person perspective images.
-Huang et al.
-(2024) built a bridge between third-person and first-person perspectives at the action level, while Dou et al.
-(2024) proposed a method designed to transform exocentric video-language data for egocentric video representation learning.
-However, research on first-person perspectives in the field of natural language processing remains unexplored.
-Figure 2: The foundation, inspiration, and detailed methods for converting the third-person ToM benchmark into a first-person perspective.
-LLMs for Interactive Scenarios
+Previous evaluations for the ToM of LLMs primarily focus on testing models using narrative stories, also referred to as reading comprehension scenarios.
+Specifically, Le et al.
+(2019) proposed the ToMi benchmark based on the classic Sally-Anne test.
+Wu et al.
+(2023) introduced the HI-ToM benchmark, which focuses on higher-order belief reasoning and sets up scenarios where agents can communicate with each other.
+Gandhi et al.
+(2023) proposed BigToM, which presents a framework for designing a ToM benchmark from synthetic templates for evaluating different aspects of LLMs’ ToM capabilities.
+Xu et al.
+(2024) introduced OpenToM, which assigns personalities to agents in the stories and ensures that the storylines are more reasonable and logical.
+Chen et al.
+(2024b) proposed ToMBench, which systematically evaluates LLMs across all dimensions of ToM capabilities.
+Unlike the above methods that require LLMs to read stories and answer related questions, some studies evaluate LLMs’ performance by inputting dialogues to them.
+Kim et al.
+(2023) proposed FanToM, which tests LLMs on their ability to infer the mental states of characters in everyday conversations.
+Chan et al.
+(2024) introduced NegotiationToM, which restricts the dialogue content to negotiation scenarios.
+For the study of LLMs’ behaviors and interaction capabilities, (Agapiou et al., 2022) proposed Melting 2.0, which encompasses various environments such as cooperation and gaming, originally designed for research in multi-agent reinforcement learning.
+(Zhou et al., 2023) introduced an interactive dialogue environment for large language models under a social goal-driven framework.
+(Chen et al., 2024a) proposed a game-like environment where different LLMs are paired for competitive interactions.
+Strategy Enhancement in Interactive Scenarios
 Some work focuses on designing interaction strategies to enable LLMs to gain more benefits during interactions.
 For example, Zhang et al.
-(2024a) proposed Agent-pro, Zhang et al.
-(2024b) introduced K-level reasoning, and Guo et al.
+(2024b) proposed Agent-pro, Zhang et al.
+(2024c) introduced K-level reasoning, and Guo et al.
 (2023) put forward the Suspicion-Agent.
@@
 (2024) examined the capability of LLMs to make rational decisions in game theoretic scenarios.
-3 EgoSocialArena
-EgoSocialArena provides a systematic approach to convert third-person perspective ToM benchmarks into first-person perspective, constructs Rule-based Agents at different cognitive levels and Reinforcement Learning (RL) Agents, and includes various scenarios to evaluate LLMs’ ToM and socialization capabilities from a first-person perspective.
-3.1 Converting Existing Third-person ToM Benchmarks to a First-person Perspective
-Foundation and Inspiration
-In LLM-based Agent applications, system message serves as a critical component, functioning to pre-set the model’s role and background.
-As illustrated in Figure 2(A), system message ”You are name and live in a town…” is used.
-Interestingly, in the domain of LLM self-awareness research (Laine et al., 2024), a similar linguistic construct is employed.
-As illustrated in Figure 2(B), researchers employ the pronoun ”you” to probe LLMs’ potential self-awareness.
-Inspired by and building upon studies in these two domains, we systematically modify system message, story, question, and answer options to transform third-person ToM benchmarks into a first-person perspective.
-Conversion Method
-As illustrated in Figure 2(C), unlike instructing LLMs in system message that ”you are a helpful assistant.”, we inform LLMs in system message that they have personally experienced certain social events, similar to deploy LLM-based Agent.
-As illustrated in Figure 2(D), we employ the pronoun ”you” to replace specific characters in stories and questions, thereby situating LLMs within particular roles.
-This approach enables the models to experience social events from a first-person perspective.
-The framing of questions is akin to that employed in self-awareness research.
-For modifications to answer options, consider LLMs answer from a first-person perspective, substituting ’I’ for specific character in the options.
-Higher-order ToM questions from third-person perspective benchmarks, after conversion, are still used to evaluate the ToM capabilities of LLMs from a first-person perspective.
-In contrast, first-order ToM questions, after conversion, form an assessment of LLMs’ mental states following social events, which we include in the scope of evaluating socialization capabilities.
-3.2 Intriguing and Distinctive Social Situations
-As illustrated in Figure 3, we design three particularly interesting scenarios—Counterfactual, New World, and Blackjack—all used to evaluate the socialization capabilities of LLMs.
-Counterfactual
-In real social situations, the rules of Rock-Paper-Scissors are: rock beats scissors, scissors beat paper, and paper beats rock.
-An LLM can relatively easily establish a belief based on this situation.
-In contrast, we define a counterfactual situation for the Rock-Paper-Scissors game (scissors beat rock, paper beats scissors, and rock beats paper) to explore whether an LLM can establish a belief that matches this counterfactual situation.
-Figure 3: Evaluation examples for the Counterfactual, New World, and Blackjack scenarios in intriguing and distinctive social situations, as well as Number Guessing scenarios in the interactive environment.
-New World
-We design stories that describe new social world scenarios that are significantly different from the current social world.
-We aim to investigate whether LLMs can demonstrate cognitive flexibility and achieve substantive shifts in their understanding and reasoning about the social world.
-Blackjack
-In the single-turn card game scenario, it is necessary to analyze the game situation, i.e., analyze own cards in Self-belief and the opponent’s cards in External-belief.
-This forms a comprehensive mental estimation of the current game situation.
-We experiment with the blackjack card game, and its specific rules can be found in A.1.
-3.3 Interactive Environment
-In the interactive environment, many studies adopt two different LLMs to interact (e.g., GPT-3.5 vs GPT-4).
-However, we have observed a phenomenon in such experiments: ”Babysitting” weaker model distracts the stronger model.
-When comparing the ToM capabilities of the Deepseek-v2 model and the GPT-3.5 model from a first-person perspective, it is unreasonable to conduct experiments like GPT-3.5 vs GPT-4 and Deepseek-v2 vs GPT-4.
-Although both opponents are GPT-4, the babysitting phenomenon during the interaction process significantly interferes with the opponent’s output, leading to unfair comparisons.
-To ensure a fair and reasonable comparison, the opponent’s capability and behavioral strategies must be controlled to be stable during the interaction process.
-We propose two methods: constructing rule-based agents at different cognitive levels as opponents and training RL agents as opponents.
-3.3.1 Rule-based Agents at Different Cognitive Levels
-Agents’ actions at lower cognitive levels follow relatively simple and fixed rules.
-As the cognitive level increases, agents’ actions adhere to more complex rule patterns, exhibiting capabilities and behavior strategies that approximate human cognitive models.
-We establish rule-based agents at different cognitive levels as opponents and denote the action of LLM Agent and rule-based Agent as $a_{m}^{t}$ and $a_{o}^{t}$ in round $t$, respectively.
-Scenario: Number Guessing
-Level 1: $a_{o}^{t}=C$.
-In this pattern, we conduct experiments with the rule-based Agent’s actions remaining constant at 50.
-Level 2: $a_{o}^{t}=f(t)=50-5(t-1)$.
-In this pattern, we conduct experiments with the rule-based Agent’s action sequence of round 1: 50, round 2: 45, …, round 9: 10, round 10: 5, an arithmetic sequence with the first term 50 and a common difference of 5.
-Level 3: $a_{o}^{t}=f(a_{m}^{t-1},a_{o}^{t-1})=0.8\times\left(\frac{a_{m}^{t-1}+a_{o}^{t-1}}{2}\right)$.
-In this pattern, we conduct experiments with the rule-based Agent’s action copying the gold value from the previous round.
-The rules of number guessing can be found in A.1.
-3.3.2 Reinforcement Learning Agents
-In the Limit Texas Hold’em scenario, we train two reinforcement learning agents as opponents: Deep Q-network (DQN)-Aggressive (Mnih et al., 2015) and DQN-Conservative (Mnih et al., 2015).
-By adapting the reward function, RL agents are given different game personalities.
-For DQN-Aggressive, we encourage the action of raising and calling during the game.
-In contrast, for DQN-Conservative, we encourage the action of folding during the game.
-The rules of Limit Texas Hold’em can be found in A.1, and a specific example of the Limit Texas Hold’em scenario can be found in A.2.
-3.3.3 Interaction
-During the interaction between an LLM and a rule-based agent in Number Guessing scenario, as illustrated in Figure 3, each round will involve asking the LLM $\mathcal{M}$ ToM questions $\mathcal{Q}_{ng}$ regarding the opponent’s behavior:
-$\mathcal{A}_{ng}^{t}=\mathcal{M}(\mathcal{Q}_{ng}^{t}\oplus\mathcal{H}^{\textless t}\oplus\mathcal{P}_{ng}),t\in\{1,2,\dots,N\}$ (1)
-where $\mathcal{P}_{ng}$ is a prompt for leading to the answer, $t$ represents the current round number, <$\mathcal{H}^{\textless t}$ represents game history information, $\mathcal{A}_{ng}^{t}$ represents LLM’s response.
-After multiple rounds of game interaction between the LLM and the RL Agent in Limit Texas Hold’em scenario, asking the LLM $\mathcal{M}$ ToM questions $\mathcal{Q}_{lth}$ regarding the opponent’s play style:
-$\mathcal{A}_{lth}^{t}=\mathcal{M}(\mathcal{Q}_{lth}^{t}\oplus\mathcal{H}^{\textless t}\oplus\mathcal{P}_{lth}),t=N$ (2)
-where $\mathcal{P}_{lth}$ is a prompt for leading to the answer, $t$ represents the current round number, <$\mathcal{H}^{\textless t}$ represents game history information, $\mathcal{A}_{lth}^{t}$ represents LLM’s response.
-4 Data construction
-4.1 Data Collection and Validation
-The conversion of the third-person perspective ToM benchmark to the first-person perspective is achieved through GPT-4o, followed by manual verification and correction.
-The game hands for Limit Texas Hold’em and Blackjack card games are generated by RLcard (Zha et al., 2019).
-Additionally, we manually construct scenarios for both the new world and counterfactual situations.
-After the data collection, following Chen et al.
-(2024)’s method, we conduct two rounds of validation to ensure the data’s correctness and quality.
-In 1st round, author A would first complete all samples created by author B.
-For stories, questions, and answer options where there are disagreements, authors A and B would discuss and modify them to reach a consensus as much as possible.
-In 2nd round, for samples where consensus is still not reached, another author C would discuss with authors A and B to determine the final answer.
-After two rounds of discussion, the final average agreement reaches 97.6%.
-4.2 Data Statistics
-EgosocialArena includes two evaluation environments: static environment and interactive environment, with seven scenarios: Daily Life, Counterfactual, New World, Blackjack, Number-Guessing, and Limit Texas Hold’em, totaling 2,195 data entries.
-A comparison with existing ToM benchmarks is shown in Table 1.
-ToM Benchmark Data Volume Evaluation Environment Number of Scenario Perspective
-ToMi 999 Static 1 Third-person
-HI-ToM 1200 Static 2 Third-person
-FanToM 254 Static 2 Third-person
-OpenToM 596 Static 3 Third-person
-ToMBench 2860 Static 1 Third-person
-EgosocialArena (Ours) 2195 Static and Dynamic 7 First-person
-Table 1: Comparison of EgosocialArena with existing ToM benchmarks.
-5 Experiments
-5.1 Experimental Setup
-We evaluate a total of 9 popular LLMs, including GPT-4o [^1] https://openai.com/index/hello-gpt-4o/, o1-preview [^2] https://openai.com/index/learning-to-reason-with-llms/, GPT-4-Turbo (Achiam et al., 2023), GPT-3.5-Turbo (Achiam et al., 2023), Claude-3.5-sonnet-20240620333https://www.anthropic.com/news/claude-3-5-sonnet, LLaMa-3-8B-Chat [^4] https://ai.meta.com/blog/meta-llama-3/, LLaMa-3-70B-Chat, LLaMa-3-8B-Instruct-Turbo, and LLaMa-3.1-405B-instruct-Turbo (Dubey et al., 2024).
-To account for the potential influence of model parameters and instruction tuning, we specifically compare LLaMa-3-8B-Chat with LLaMa-3-8B-Instruct-Turbo, as well as LLaMa-3-8B-Chat with LLaMa-3-70B-Chat.
-To establish a human performance baseline, we recruit 10 graduate students, all of whom have received a good basic education and possess mature cognitive abilities, to complete responses to the questions in EgoSocialArena.
-The average accuracy of their responses will serve as the human performance baseline.
-No extra tutorials or examples are provided to ensure a fair comparison.
-5.2 Evaluation Method
-For Daily Life, New World, and CounterFactful scenarios, we present LLMs with a story, a question, and several options, then ask them to pick the correct answer.
-Using the accuracy of answering questions as the evaluation metric for these scenarios.
-In the interactive environments of Number Guessing and Texas Hold’em, the evaluation of these scenarios also has standard answers because we propose agents with stable capabilities and behavioral strategies as opponents of LLMs.
-For the Blackjack scenario, we conducted a manual evaluation.
-To ensure the quality of the manual evaluation, we measure the average consistency score between evaluators, which reached 96.3%.
-Methods ToMI Number Guessing Texas Hold’em
-Third-person First-person Level 1 Level 2 Level 3
-Open-source Models
-LLaMa-3-8B-Chat 50.6 66.2 0.0 0.0 0.0 48.0
-LLaMa-3-70B-Chat 58.4 63.2 10.0 20.0 10.0 38.0
-LLaMa-3-8B-Instruct 51.1 65.4 0.0 0.0 0.0 44.0
-LLaMa-3.1-405B-Instruct 58.0 65.8 80.0 20.0 20.0 56.0
-API-based Models
-Claude-3-5-Sonnet 71.0 80.5 50.0 10.0 40.0 66.0
-GPT-3.5-Turbo 45.5 51.9 10.0 10.0 0.0 56.0
-GPT-4-Turbo 55.4 69.7 10.0 20.0 10.0 60.0
-GPT-4o 64.1 71.0 10.0 40.0 10.0 62.0
-o1-preview 71.9 77.5 90.0 90.0 90.0 72.0
-Human
-Human Performance 90.2 90.2 90.0 86.0 73.0 82.0
-Table 2: The performance of ToM capabilities from the first-person perspective in open-source models and API-based models.
-The highest and second-highest scores among models and humans in each section are highlighted in blue and red, respectively.
-5.3 Experimental Results
-The performance of ToM and socialization capabilities from the LLMs’ first-person perspective is shown in Table 2 and 3, respectively.
-Based on the experimental results, we have obtained some key insights:
-Performance Differences in LLMs’ Theory of Mind (ToM) Across Third-Person and First-Person Perspective
-As shown in Table 2, all LLMs exhibited improved performance after the ToMI dataset was adapted from a third-person to a first-person perspective.
-Notably, the Claude and o1-preview models demonstrated significantly stronger ToM capabilities in the first-person perspective compared to other models.
-Except for GPT-3.5-Turbo, API-based models generally outperformed open-source models, including the recently released LLaMa-3.1-405B-Instruct.
-However, despite these improvements, there remains a substantial gap between the performance of all LLMs and that of human participants.
-Methods New World Counterfact Blackjack ToMBench
-All Desire Emotion Intention
-Open-source Models
-LLaMa-3-8B-Chat 6.7 71.0 84.5 67.4 71.5 73.0 50.0
-LLaMa-3-70B-Chat 13.3 59.0 92.5 73.5 72.5 63.5 83.0
-LLaMa-3-8B-Instruct 6.7 64.0 81.4 55.1 45.8 70.5 53.5
-LLaMa-3.1-405B-Instruct 36.7 66.0 96.5 77.6 75.0 75.5 80.6
-API-based Models
-Claude-3-5-Sonnet 90.0 74.0 96.9 79.6 72.0 64.5 82.6
-GPT-3.5-Turbo 13.3 37.0 89.7 72.5 70.0 86.0 68.5
-GPT-4-Turbo 23.3 70.0 95.0 75.6 84.5 67.4 84.0
-GPT-4o 36.7 52.0 96.5 85.7 78.0 91.0 96.5
-o1-preview 86.7 90.0 97.2 84.8 73.0 91.0 98.0
-Human
-Human Performance 93.3 91.0 97.0 90.7 83.2 92.0 97.5
-Table 3: The performance of socialization capabilities from the first-person perspective of open-source models and API-based models (after the ToMBench conversion, it also includes part of the ToM capability assessment, which we have consolidated here together) with the highest and second-highest scores among models and humans in each section are highlighted in blue and red, respectively.
-The powerful capabilities of the o1-preview model are truly surprising
-In the Number-Guessing scenario, almost all large language models perform poorly, even in the simplest Level 1 situation, which poses a significant challenge for humans as well.
-However, the recent o1-preview model has performed exceptionally well, demonstrating a cognitive level in this scenario that approaches or even surpasses that of humans.
-The scaling up of open-source models or performing instruction fine-tuning has not yielded significant results
-By comparing the performance of LLaMa-3-8B-Chat with LLaMa-3-70B-Chat, as well as LLaMa-3-8B-Chat with LLaMa-3-8B-Instruct models in Table 2, 3, it is observed that neither of these measures has notably improved the ToM and socialization capabilities of LLMs.
-To enhance LLMs’ ToM and socialization capabilities, innovative approaches will be needed in future research on LLMs.
-The performance gap between open-source models and API-based models in socialization scenarios is significant
-In counterfactual and New World scenarios, API-based models exhibit extremely flexible cognition, demonstrating a strong ability to adapt to social situations and showing performance comparable to humans.
-In contrast, open-source models perform poorly.
-For instance, the LLaMa-3-8B-Chat model only achieves an accuracy rate of 6.7% in New World scenarios, compared to 90.0% for the Claude model.
-In the dimensions of desire and intention in social contexts, LLMs have shown potential
-As illustrated in Table 3, although LLMs still lag behind humans in the overall dimension of mental states, their performance in the desire and intention dimensions is impressive.
-These findings can further inspire and serve as a foundation for the application of LLM-based agents in the real social world.
-Figure 4: In the scenario of Number Guessing Level 2 (Arithmetic sequence), the belief state evolution pattern of GPT-4-Turbo regarding the opponent’s proposed numbers.
-5.4 Qualitative Analysis
-Analysis of Model Failure Causes and Behavioral Patterns in Number Guessing Scenario —— Mid-point Belief, Strange Guess and Get Back on Track
-As shown in Figure 4, in the scenario of Number Guessing (Level 2: Arithmetic sequence), we thoroughly investigate the belief state evolution pattern of GPT-4-Turbo regarding the opponent’s proposed numbers.
-In round 1, with no available information, the GPT-4-Turbo model thinks the opponent will choose the number 50 within the range of 1-100.
-The same phenomenon is observed in the GPT-3.5-Turbo model, called ”mid-point belief”.
-Sometimes, the GPT-4-Turbo model continuously believes the opponent will choose progressively smaller numbers throughout the entire interaction, as depicted by the GPT-4-Turbo guess1 curve in Figure 4.
-Although this is very close to the gold number, it does not capture that the opponent’s chosen numbers form an arithmetic sequence.
-Another situation occurs when the GPT-4-Turbo model makes a ”strange guess” in the initial rounds, thinking the opponent will suddenly choose larger numbers.
-After several rounds, it captures that the opponent’s chosen numbers form an arithmetic sequence, called ”Get Back on Track”.
-Overall, despite the statistical results indicating that the GPT-4-Turbo model does not establish a belief regarding the Level 2 opponent in the Number Guessing scenario, the phenomena we observed suggest that it has started to grasp some patterns.
-Analysis of our research and current studies centered on outcomes and benefits
-Many current works focus on designing various strategies to improve the performance of LLMs in static or interactive environments.
-For example, making an LLM display anger to enhance its performance in negotiation scenarios, or using carefully crafted prompt engineering to analyze an opponent’s behavior from multiple perspectives to achieve higher payoffs.
-Unlike these approaches, we fundamentally explore the ToM capabilities of LLMs from a first-person perspective in various scenarios.
-Based on the behavior patterns of LLMs that we observed in our experiments, such as the excellent performance of the o1-preview and Claude models in the situated new world scenarios, it indicates that these models have significant potential for further exploration in fields like role-playing and simulation, as well as entering the real social world.
-6 Conclusion
-In this paper, considering the limitations of existing ToM and socialization benchmarks, the importance of first-person ToM and socialization capabilities in LLMs, and the natural approach of observing and understanding the world from an ego-centric first-person perspective for both humans and AI agents, we propose the EgoSocialArena framework.
-This framework is designed to comprehensively evaluate and probe the first-person ToM and socialization capabilities of LLMs in both static and interactive environments, covering multiple scenarios.
-To avoid the ”babysitting” problem during the evaluation process and achieve fair and comprehensive assessments, we construct rule-based agents at different cognitive levels and train RL agents.
-We collect a total of 2195 test data entries and test multiple advanced and popular LLMs, revealing some interesting and key insights, and highlighting a significant potential for enhancing the first-person ToM and socialization capabilities of LLMs, allowing them to truly enter the social world.
-Appendix A Appendix
-A.1 Game Rules
-Blackjack
-Blackjack, also known as 21, is a popular card game that involves a dealer and a player.
-Players must decide whether to hit or stand based on their own hand, the dealer’s face-up card, and the dealer’s one hidden card.
-The objective is to beat the dealer without exceeding 21 points.
-Limit Texas Hold’em
-The game commences with each player being dealt two private cards, which belong exclusively to the player and remain hidden from the others.
-Five community cards are then dealt face-up in a series of stages: a three-card Flop, followed by a single card on the Turn and another single card on the River.
-The player can choose from four actions: Fold, Check, Call, Raise.
-They aim to construct the best five-card poker hand possible using any combination of their private cards and community cards.
-Number Guessing
-Each player selects a number between 1 to 100.
-The objective is to select a number that is closest to 80% of the group’s average choice.
-The key idea is to guess how others will estimate the average and thus decide the number to submit.
-A.2 Limit Texas Hold’em Scenario
-Figure 5: Evaluation examples for the Limit Texas Hold’em Scenario.
+Zhang et al.
+(2024a) propose to optimize the structure of thought.
+Necessity of developing LLMs’ Social Intelligence
+With LLMs becoming increasingly integrated into our everyday lives, developing LLMs with social intelligence could be better at communicating with us, collaborating with us, understanding us, teaching us and learning from us.
+(Gandhi et al., 2021, 2023; Rabinowitz et al., 2018; Shu et al., 2021).
+Appendix B Case——Limit Texas Hold’em
+As illustrated in Figure 6.
+Figure 6: A Case for Limit Texas Hold’em.
+Appendix C Belief Dynamic Evolution in G0.8A Scenario
+The following three tables correspond to the dynamic evolution data of beliefs for various LLMs under Cognitive Levels 3, 2, and 1.
+Model Round 1 Round 2 Round 3 Round 4 Round 5 Round 6 Round 7 Round 8 Round 9 Round 10 Accuracy
+GPT-4-Turbo 50 45 40 35 30 25 22 17 15 13 0.1
+GPT-3.5-Turbo 40 20 60 55 70 90 60 45 75 85 0
+GPT-4o 50 40 30 20 15 10 8 6 5 4 0.1
+o1-preview 1 20 0.9
+Claude-3-5-Sonnet-20240620 65 45 35 28 20 17 14 10 7.5 5.6 0.4
+Llama3-8b-chat-hf 67 67 67 67 67 67 67 67 67 67 0
+Llama3-70b-chat-hf 50 45 43 30 25 19 15 12 11 7 0.1
+Llama3.1-405b-Instruct-Turbo 50 40 35 29 23 19 14.5 11.5 9.5 7.5 0.2
+Model Round 1 Round 2 Round 3 Round 4 Round 5 Round 6 Round 7 Round 8 Round 9 Round 10 Accuracy
+GPT-4-Turbo 50 45 48 42 36 33 28 22 18 12 0.2
+GPT-3.5-Turbo 40 20 60 35 70 50 45 60 45 40 0.1
+GPT-4o 50 40 40 30 25 20 15 10 10 5 0.4
+o1-preview 1 0.9
+Claude-3-5-Sonnet-20240620 65 45 35 25 20 15 12 8 5 8 0.1
+Llama3-8b-chat-hf 67 67 67 67 67 67 67 67 67 67 0
+Llama3-70b-chat-hf 50 45 38 32 28 24 21 19 16 11 0.1
+Llama3.1-405b-Instruct-Turbo 50 40 35 30 28 25 22 18 15 10 0.2
+Model Round 1 Round 2 Round 3 Round 4 Round 5 Round 6 Round 7 Round 8 Round 9 Round 10 Accuracy
+GPT-4-Turbo 50 45 48 47 48 49 48 47 46 45 0.1
+GPT-3.5-Turbo 40 35 70 30 80 40 55 60 50 30 0.1
+GPT-4o 50 40 30 40 35 45 45 45 45 45 0.1
+o1-preview 1 0.9
+Claude-3-5-Sonnet-20240620 65 45 35 25 20 50 50 50 50 50 0.5
+Llama3-8b-chat-hf 67 67 67 67 67 67 67 67 67 67 0
+Llama3-70b-chat-hf 50 48 52 53 54 55 54 56 57 58 0.1
+Llama3.1-405b-Instruct-Turbo 50 33 45 50 50 50 50 50 50 50 0.8

### v3

@@
-Entering Real Social World! Benchmarking the Social Intelligence of Large Language Models from a First-person Perspective
+EgoSocialArena: Benchmarking the Social Intelligence of Large Language Models from a First-person Perspective
 Abstract
-Social intelligence is built upon three foundational pillars: cognitive intelligence, situational intelligence, and behavioral intelligence.
-As large language models (LLMs) become increasingly integrated into our social lives, understanding, evaluating, and developing their social intelligence are becoming increasingly important.
+Social intelligence is built upon three foundational pillars: cognitive, situational, and behavioral intelligence.
+As large language models (LLMs) are increasingly integrated into our social lives, understanding, evaluating, and developing their social intelligence are becoming increasingly important.
 While multiple existing works have investigated the social intelligence of LLMs, (1) most focus on a specific aspect, and the social intelligence of LLMs has yet to be systematically organized and studied; (2) position LLMs as passive observers from a third-person perspective, such as in Theory of Mind (ToM) tests.
@@
 In light of this, we present EgoSocialArena, a novel framework grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence, aimed to systematically evaluate the social intelligence of LLMs from a first-person perspective.
-With EgoSocialArena, we conduct a comprehensive evaluation of eight prominent foundation models, even the most advanced LLMs like o1-preview lag behind human performance by 11.0 points.
-Our code and data are available at https://github.com/gyhou123/EgoSocialArena.
+Using EgoSocialArena, we conduct a comprehensive evaluation of eight prominent foundation models.
+Our findings show that even the most advanced LLMs, such as O1-preview, still fall significantly behind human performance [^1] https://github.com/gyhou123/EgoSocialArena.
 1 Introduction
@@
 The development of EgoSocialArena is grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence: (1) Cognitive Intelligence: we propose a complete and generalizable workflow to transform existing static third-person ToM benchmarks into a first-person perspective.
-Additionally, we have newly developed a dynamic cognitive assessment in multi-turn interactive scenarios.
+Additionally, by constructing rule-based agents and reinforcement learning agents with stable capability levels and behavior strategies, we have newly developed a dynamic cognitive assessment in multi-turn interactive scenarios.
 (2) Situational Intelligence: Imagine an LLM-based Agent entering our social world - how would it respond emotionally when receiving praise or gifts [^2] This might be related to self-awareness, but the focus could be shifted more towards the application situations.?
@@
 We establish a human performance baseline by engaging qualified human annotators with a college degree or higher.
-Our experimental results reveal several interesting and critical insights: (1) The o1-preview model achieved the highest score of 80.6 among all models, surpassing human performance in dynamic cognition and adversarial game scenarios.
-Nevertheless, an 11.0 gap in overall accuracy remains when compared to the human baseline, leaving plenty of room for model improvement.
-Our in-depth analysis reveals that the superiority of o1-preview is mainly attributed to its powerful logical reasoning and mathematical abilities (keenly uncovering deeper patterns behind the data).
-(2) Comparing the performance of LLaMA3-8B-Chat and LLaMA3-70B-Chat models shows that simply scaling model size does not significantly help improve the social intelligence of LLMs.
+Our experimental results reveal several interesting and critical insights: (1) The o1-preview model achieved the highest score of 80.6, surpassing human performance in dynamic cognition and adversarial game scenarios.
+Nevertheless, an 7.7 gap in overall accuracy remains when compared to the human baseline, leaving room for improvement.
+Our analysis reveals that the superiority of o1-preview is mainly attributed to its powerful logical reasoning and mathematical abilities (uncovering deeper patterns behind the data).
+(2) Comparing the performance of LLaMA3-8B and LLaMA3-70B models show that scaling model size does not significantly improve the social intelligence of LLMs.
 (3) Compared to the third-person perspective, LLMs show significantly improved ToM reasoning ability when operating from a first-person perspective.
@@
 4.1 Experimental Setup
-We evaluate a total of eight prominent foundation LLMs, including GPT-4o [^2] https://openai.com/index/hello-gpt-4o/, o1-preview [^3] https://openai.com/index/learning-to-reason-with-llms/, GPT-4-Turbo (Achiam et al., 2023), GPT-3.5-Turbo (Achiam et al., 2023), Claude-3.5-sonnet-20240620444https://www.anthropic.com/news/claude-3-5-sonnet, LLaMa-3-8B-Chat [^5] https://ai.meta.com/blog/meta-llama-3/, LLaMa-3-70B-Chat, and LLaMa-3.1-405B-instruct-Turbo (Dubey et al., 2024).
+We evaluate a total of eight prominent foundation LLMs, including GPT-4o [^3] https://openai.com/index/hello-gpt-4o/, o1-preview [^4] https://openai.com/index/learning-to-reason-with-llms/, GPT-4-Turbo (Achiam et al., 2023), GPT-3.5-Turbo (Achiam et al., 2023), Claude-3.5-sonnet-20240620555https://www.anthropic.com/news/claude-3-5-sonnet, LLaMa-3-8B-Chat [^6] https://ai.meta.com/blog/meta-llama-3/, LLaMa-3-70B-Chat, and LLaMa-3.1-405B-instruct-Turbo (Dubey et al., 2024).
 To account for the potential influence of model parameters, we specifically compare LLaMa-3-8B-Chat with LLaMa-3-70B-Chat.
-To establish a human performance baseline, we recruit 10 graduate students, all of whom have received a good basic education and possess mature cognitive abilities, to complete responses to the questions in EgoSocialArena.
+To establish a reliable human performance baseline, we recruit 50 graduate students, all of whom have received a good education and possess excellent cognitive abilities, to complete responses to the questions in EgoSocialArena.
 The average accuracy of their responses will serve as the human performance baseline.
@@
 For the Social-goal driven interactive dialogue, we use GPT-4 to automatically evaluate the performance of humans and LLMs in terms of goal completion during their interactions.
-Methods Cognitive Intelligence Static Cognition Dynamic Cognition-G0.8A Dynamic Cogntion Third-person First-person $\Delta$ Level 1 Level 2 Level 3 Limit Texas Open-source Models LLaMa-3-8B-Chat 50.6 66.2 +15.6 0.0 0.0 0.0 48.0 LLaMa-3-70B-Chat 58.4 63.2 +4.8 10.0 20.0 10.0 38.0 LLaMa-3.1-405B-Instruct 58.0 65.8 +7.8 80.0 20.0 20.0 56.0 API-based Models Claude-3-5-Sonnet 71.0 80.5 +9.5 50.0 10.0 40.0 66.0 GPT-3.5-Turbo 45.5 51.9 +6.4 10.0 10.0 0.0 56.0 GPT-4-Turbo 55.4 69.7 +14.3 10.0 20.0 10.0 60.0 GPT-4o 64.1 71.0 +6.9 10.0 40.0 10.0 62.0 o1-preview 71.9 77.5 +5.6 90.0 90.0 90.0 72.0 Human Human Performance 90.2 90.2 0.0 90.0 86.0 73.0 82.0
-Methods Situational Intelligence Behavioral Intelligence AVG Parallel World Counterfact Real-World Adversarial Cooperative Social Goal Open-source Models LLaMa-3-8B-Chat 6.7 71.0 67.2 51.3 49.7 22.5 34.8 LLaMa-3-70B-Chat 13.3 59.0 73.2 45.0 53.3 25.5 37.3 LLaMa-3.1-405B-Instruct 36.7 66.0 77.3 52.3 65.2 34.0 52.1 API-based Models Claude-3-5-Sonnet 90.0 74.0 79.8 55.0 94.8 50.5 62.8 GPT-3.5-Turbo 13.3 37.0 72.2 46.7 50.3 33.0 34.6 GPT-4-Turbo 23.3 70.0 75.7 54.7 75.6 52.0 47.4 GPT-4o 36.7 52.0 85.8 54.0 80.8 53.0 50.5 o1-preview 86.7 90.0 84.7 56.7 96.3 52.5 80.6 Human Human Performance 93.3 91.0 90.7 56.3 100.0 64.5 91.6
+Methods Cognitive Intelligence Static Cognition Dynamic Cognition-G0.8A Dynamic Cogntion Third-person First-person $\Delta$ Level 1 Level 2 Level 3 Limit Texas Open-source Models LLaMa-3-8B-Chat 50.6 66.2 +15.6 0.0 0.0 0.0 48.0 LLaMa-3-70B-Chat 58.4 63.2 +4.8 10.0 20.0 10.0 38.0 LLaMa-3.1-405B-Instruct 58.0 65.8 +7.8 80.0 20.0 20.0 56.0 API-based Models Claude-3-5-Sonnet 71.0 80.5 +9.5 50.0 10.0 40.0 66.0 GPT-3.5-Turbo 45.5 51.9 +6.4 10.0 10.0 0.0 56.0 GPT-4-Turbo 55.4 69.7 +14.3 10.0 20.0 10.0 60.0 GPT-4o 64.1 71.0 +6.9 10.0 40.0 10.0 62.0 o1-preview 71.9 77.5 +5.6 90.0 90.0 90.0 72.0 Human Human Performance 97.4 97.4 0.0 90.0 89.0 85.0 94.0
+Methods Situational Intelligence Behavioral Intelligence AVG Parallel World Counterfact Real-World Adversarial Cooperative Social Goal Open-source Models LLaMa-3-8B-Chat 6.7 71.0 67.2 51.3 49.7 22.5 34.8 LLaMa-3-70B-Chat 13.3 59.0 73.2 45.0 53.3 25.5 37.3 LLaMa-3.1-405B-Instruct 36.7 66.0 77.3 52.3 65.2 34.0 52.1 API-based Models Claude-3-5-Sonnet 90.0 74.0 79.8 55.0 94.8 50.5 62.8 GPT-3.5-Turbo 13.3 37.0 72.2 46.7 50.3 33.0 34.6 GPT-4-Turbo 23.3 70.0 75.7 54.7 75.6 52.0 47.4 GPT-4o 36.7 52.0 85.8 54.0 80.8 53.0 50.5 o1-preview 86.7 90.0 84.7 56.7 96.3 52.5 80.6 Human Human Performance 96.7 97.0 96.3 56.6 100.0 69.0 88.3
 Table 2: Performance of cognitive, situational, and behavioral intelligence from first-person perspective of eight LLMs.
@@
 As shown in Table 2, the o1-preview model achieved the highest score of 80.6 among all models, surpassing human performance in dynamic cognition and adversarial game scenarios.
-Nevertheless, an 11.0 gap in overall performance remains when compared to the human baseline, leaving plenty of room for model improvement.
+Nevertheless, an 7.7 gap in overall performance remains when compared to the human baseline, leaving plenty of room for model improvement.
 The second-best performer is the claude-3-5-sonnet model, which demonstrate impressive results in the static cognition and parallel world scenarios.
@@
 Overall, the performance of open-source models lags significantly behind that of API-based models and most models still exhibit a large performance gap compared to humans.
-For instance, the LLaMa-3-8B-Chat model achieved an overall score of 34.8, significantly lower than the human performance of 91.6.
+For instance, the LLaMa-3-8B-Chat model achieved an overall score of 34.8, significantly lower than the human performance of 88.3.
 4.4 In-Depth Analysis
@@
 5 Conclusion
-In this paper, considering the social intelligence of LLMs has yet to be systematically organized and studied, ego-centric first-person perspective evaluation can align well with actual LLM-based Agent use scenarios, incorporating human-machine interaction scenario evaluation is critical and the natural approach of observing and understanding the world from an ego-centric first-person perspective for both humans and LLM-based agents, we propose the EgoToMArena framework.
-This framework is grounded in the three pillars of social intelligence: cognitive, situational, and behavioral intelligence, with eight scenarios: static, dynamic cognition; real-world, counterfactual, parallel world situation; cooperative, adversarial game, and social goal-driven human-machine interactive environment, aimed to systematically evaluate the social intelligence of LLMs from a first-person perspective.
-We conduct extensive experiments and observe some key insights regarding the future development of LLMs as well as the capabilities levels of the most advanced LLMs currently available.
+In this paper, we propose EgoSocialArena, a novel framework grounded in the three pillars of social intelligence: cognitive intelligence, situational intelligence, and behavioral intelligence, designed to systematically evaluate the social intelligence of LLMs from a first-person perspective.
+EgoSocialArena incorporates several unique design elements, including third-person to first-person perspective transformation, constructing rule-based agents and reinforcement learning agents with stable capabilities levels and behavior strategies for dynamic cognition assessment, considering non-standard and atypical social situations, evaluating the mental states of LLMs’ self after experiencing certain social events (this may be related to self-awareness), and exploring human-machine interaction.
+We conduct comprehensive experiments and observe some key insights regarding the future development of LLMs as well as the capabilities levels of the most advanced LLMs currently available.
 Limitations